# A Simple, Optimal and Efficient Algorithm for Online Exp-Concave Optimization

### v1

@@ A.2 Proof of Lemma˜1
Since $g(\mathbf{x})=\mathrm{e}^{-\alpha f(\mathbf{x})}$ is concave and positive, then $h(\mathbf{x})=\mathrm{e}^{-2\gamma_{0}f(\mathbf{x})}$ is also concave because $2\gamma_{0}/\alpha\leq 1$ and, for any $\lambda\in[0,1]$ and any $(\mathbf{x},\mathbf{y})\in\mathcal{X}^{2}$,
$\geq\lambda g(\mathbf{x})^{2\gamma_{0}/\alpha}+(1-\lambda)g(\mathbf{y})^{2\gamma_{0}/\alpha}=\lambda h(\mathbf{x})+(1-\lambda)h(\mathbf{y})$
where the first inequality follows from the concavity of $g$ and the second from the concavity of $a\mapsto a^{2\gamma_{0}/\alpha}$. Then, by the concavity of $h$, for any $(\mathbf{x},\mathbf{u})\in\mathcal{X}^{2}$,
Rearranging the preceding inequality yields
To prove Eq.˜3, it suffices to apply that $\log(1+a)\leq a-a^{2}/4$ (which holds for any $\left|a\right|\leq 1$) to the right-hand side of the preceding inequality. We note that $\left|2\gamma_{0}\nabla f(\mathbf{x})^{\mathsf{T}}(\mathbf{x}-\mathbf{u})\right|\leq 1$ due to the selection of $\gamma_{0}=\frac{1}{2}\min\left\{\frac{1}{DG},\alpha\right\}$. ∎
Appendix B Proofs for Section˜3
In this section, we prove the guarantees of LightONS.
B.1 Proof of Lemma˜2
Proof.
Let $\Phi_{T}$ denote the sum of squared norms of updates, as in the left-hand side of Eq.˜36. Note that $A_{t}=\epsilon I+\sum_{i=1}^{t}\nabla f_{t}(\mathbf{x}_{t})\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}$ in Algorithm˜1 and $\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}\leq G$ in Assumption˜2. Then by Lemma˜6 we have
@@ B.1 Proof of Lemma˜2
Let $\left\{\tau_{i}\right\}_{i=1}^{N}\subseteq[T]$ denote all the rounds when LightONS.Core performs the Mahalanobis projection. With $\tau_{0}=1$, we have that for any $i\in[N]$,
@@
We remark that Lemma˜2 is tight in terms of $T$ up to logarithmic factors. Consider $d=1$ and $\nabla f_{t}(\mathbf{x}_{t})=1/\sqrt{T}$, then $A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})=\Omega(1/\sqrt{T})$ and $\|\sum_{t=1}^{T}A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})\|_{2}=\Omega(\sqrt{T})$.
B.2 Proof of Theorem˜1
@@
The proof of Theorem˜1 follows as a specialization of the proof of Theorem˜2 given in Section˜B.4. The only substantive difference is that, in the present setting, the analysis is carried out directly on the original loss functions $f_{t}$, rather than on surrogate losses $g_{t}$.
@@
  • Choice of the expanded curvature parameter $\gamma$ in Algorithm˜2. The expression $\gamma=\frac{1}{2}\min\{\frac{2}{(k+1)DG},\alpha\}$ directly follows from Lemma˜1.
  • OMD regret analysis with selective projections. The regret analysis follows the same structure as Lemma˜8, which accounts for projections triggered only when the iterate exits the expanded domain $\widetilde{\mathcal{X}}$. Details are given in Section˜B.4.
  • Runtime of LightONS.Core. The runtime bound is derived in the same manner as in the proof of Theorem˜2 in Section˜B.4.
B.3 Proof of Lemma˜4
@@
By the definition of curvature parameters, namely $\gamma_{0}=\frac{1}{2}\min\left\{\frac{1}{DG},\alpha\right\}$ in Lemma˜1 and $\gamma^{\prime}=\frac{1}{2}\min\left\{\frac{1}{c_{f}c_{g}DG},\frac{4}{(k+1)c_{f}c_{g}DG},\alpha\right\}$ in Algorithm˜3, we first have $\gamma^{\prime}\leq\gamma_{0}$. Then by relaxing Eq.˜3 in Lemma˜1 we have
@@
$\leq\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})-\frac{\gamma^{\prime}}{2}\left(\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\right)^{2}$
Let $U(a)=a-\frac{\gamma^{\prime}}{2}a^{2}$. To finish the proof, it suffices to show that
@@
Note that $U$ is monotonically increasing on $(-\infty,\frac{1}{\gamma^{\prime}}]$. Therefore, it suffices to show that
@@
The first inequality follows from Condition˜1; The second inequality from the Cauchy-Schwarz inequality $\mathbf{u}^{\mathsf{T}}\mathbf{v}\leq\left\|\mathbf{u}\right\|_{2}\left\|\mathbf{v}\right\|_{2}$; The last inequality from the definition of $\gamma^{\prime}$. ∎
B.4 Proof of Theorem˜2
@@
Based on Lemma˜7, we can decompose the regret of LightONS as the following lemma shows.
@@
Proof of Lemma˜8.
When $\left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}\leq kD/2$, no Mahalanobis projection is performed, i.e., $\mathbf{y}_{t+1}=\widehat{\mathbf{y}}_{t+1}=\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})$, then for any $\mathbf{u}\in\mathbb{R}^{d}$, we have
@@
Otherwise, when $\left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}>kD/2$, a Mahalanobis projection is performed, by Lemma˜7, for any $\mathbf{u}\in\mathcal{X}$, we have
@@
$\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}\leq\left\|\left(\mathbf{y}_{t}-\mathbf{u}\right)-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})\right\|_{A_{t}}^{2}.$ (39)
Rearranging Eq.˜39 yields the desired result of Eq.˜38. ∎
Note that Lemma˜8 ignores the truncation error of the Mahalanobis projections. The next lemma complements the analysis by showing that, thanks to FastProj (Algorithm˜4), the truncation error only incurs a negligible additive $O(1/t^{2})$ term in the regret decomposition, which can be safely ignored in the final regret bound of LightONS.
@@
Let the Mahalanobis projection of Algorithm˜3 is implemented with Algorithm˜4 with $A=A_{t}$, $R=D/2$, $\mathbf{u}=\widehat{\mathbf{y}}_{t+1}$, $\zeta=\zeta_{t}$, $\underline{\lambda}=\epsilon$, $\overline{\lambda}=\epsilon+G^{2}t$, where $\zeta_{t}$ is defined as
@@
Proof of Lemma˜9.
It suffices to prove the case when $\left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}>kD/2$. Let $\mathbf{y}^{\ast}_{t+1}=\Pi_{\mathcal{B}(D/2)}^{A_{t}}\left[\widehat{\mathbf{y}}_{t+1}\right]$ denote the exact Mahalanobis projection without truncation error, and $\bm{\delta}_{t}=\mathbf{y}^{\ast}_{t+1}-\mathbf{y}_{t+1}$ denote the truncation error. By Lemma˜11, $\left\|\bm{\delta}_{t}\right\|_{2}=\left\|\mathbf{y}_{t+1}-\mathbf{y}^{\ast}_{t+1}\right\|_{2}\leq\zeta_{t}$; By Lemma˜7, for any $\mathbf{u}\in\mathcal{X}$,
@@
$\leq\left\|\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})-\mathbf{u}\right\|_{A_{t}}^{2}+\frac{2\gamma^{\prime}}{t^{2}}$
The first inequality considers the operator norm of $A_{t}$ and uses the fact that $\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}\leq c_{g}G$ in Condition˜1; The second inequality comes from the selection of $\zeta_{t}$ in Eq.˜40. ∎
@@
Substituting Eq.˜38 into Eq.˜13 yields the following inequalities:
@@
$=$ $\frac{1}{2\gamma^{\prime}}\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{A_{t}^{-1}}^{2}+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{A_{t-1}}^{2}-\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}$
Telescoping the preceding inequality establishes the desired regret bound Eq.˜14:
@@
Finally, the desired runtime Eq.˜15 follows from the following two parts:
@@
  • Runtime aside from FastProj. In each round, the domain conversion of Lemma˜3 takes $O(\texttt{EP}_{\mathcal{X}}+d)$ time, updating and inverting $A_{t}$ as Eq.˜18 takes $O(d^{2})$ time, and other operations, such as computing the surrogate gradient as Eq.˜12b, take only $O(d)$ time. Overall runtime aside from FastProj is $O((\texttt{EP}_{\mathcal{X}}+d^{2})T)$.
@@
  • Runtime of FastProj. In each call of FastProj, by Theorem˜3, the iteration number of bisections $n_{t}=O(\log t)=O(\log T)$ and each bisection takes $O(d^{\omega})$ time with Choice 1. Since LightONS.Core is a subroutine in LightONS, the number of calls to FastProj is at most $O((k-1)^{-1}d^{0.5}\sqrt{T/\epsilon})$ by Lemma˜2. Overall runtime of FastProj is $O((k-1)^{-1}d^{0.5}\sqrt{T/\epsilon}\cdot d^{\omega}\log T)$.
B.5 Proof of Theorem˜3
First we reduce the Mahalanobis projection onto an Euclidean ball to a one-dimensional root-finding problem as the following lemma shows.
Lemma 10.
$0=\rho(\mu)=\left\|\left(A+\mu I\right)^{-1}A\mathbf{y}\right\|_{2}^{2}-R^{2}=\sum_{i=1}^{d}\frac{v_{i}^{2}}{\left(1+\frac{\mu}{\lambda_{i}}\right)^{2}}-R^{2}.$ (42)
Here $v_{i}=\mathbf{e}_{i}^{\mathsf{T}}Q^{\mathsf{T}}\mathbf{y}$, where $\mathbf{e}_{i}$ is the $i$-th standard basis vector, and $A=Q\Lambda Q^{\mathsf{T}}$ is the eigendecomposition, $\Lambda=\operatorname{diag}\left(\lambda_{1},\dots,\lambda_{d}\right)$ with $\lambda_{1}\geq\dots\geq\lambda_{d}>0$.
Besides, the positive zero of $\rho$, denoted as $\mu^{\ast}$, satisfies that
$\left(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1\right)\lambda_{d}<\mu^{\ast}<\left(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1\right)\lambda_{1}.$
Proof of Lemma˜10.
The Mahalanobis projection onto an Euclidean ball is formulated as
$\min_{\mathbf{x}\in\mathbb{R}^{d}}$ $(\mathbf{x}-\mathbf{y})^{\mathsf{T}}A(\mathbf{x}-\mathbf{y})$
s.t. $\mathbf{x}^{\mathsf{T}}\mathbf{x}\leq R^{2}$
The Lagrangian of this quadratic program is $\mathcal{L}=(\mathbf{x}-\mathbf{y})^{\mathsf{T}}A(\mathbf{x}-\mathbf{y})+\mu(\mathbf{x}^{\mathsf{T}}\mathbf{x}-R^{2})$. According to the KKT conditions, $\nabla\mathcal{L}=2A(\mathbf{x}-\mathbf{y})+2\mu\mathbf{x}=\bm{0}$ with $\mu>0$. (Otherwise, $\mu=0$ implies $\mathbf{y}\in\mathcal{B}(R)$ and the projection is trivial.) Rearranging $\nabla\mathcal{L}=\bm{0}$ yields $\mathbf{x}=(A+\mu I)^{-1}A\mathbf{y}$, thus $\mathbf{x}^{\mathsf{T}}\mathbf{x}=R^{2}$ is equivalent to $\rho(\mu)=0$, which proves Eq.˜42.
Note that $\sum_{i=1}^{d}v_{i}^{2}=\left\|Q\mathbf{y}\right\|_{2}^{2}=\left\|\mathbf{y}\right\|_{2}^{2}$ due to the orthogonality of $Q$. Therefore,
It is straightforward that $\rho((\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{d})\geq 0$ and $\rho((\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{1})\leq 0$. Then, since $\rho$ monotonically decreases on $0,\infty)$, we conclude that $\mu^{\ast}\in((\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{d},(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{1})$ and $\mu^{\ast}$ is unique on $0,\infty)$. ∎
The next lemma reveals the relationship between the truncation error of the root-finding and the truncation error of the Mahalanobis projection.
Lemma 11.
Let $\mathbf{x}^{\ast}=(A+\mu^{\ast}I)^{-1}A\mathbf{y}$, where $\rho(\mu^{\ast})=0$ as in Lemma˜10. If $\widetilde{\mu}\geq 0$ and $\left|\widetilde{\mu}-\mu^{\ast}\right|<\frac{\lambda_{d}}{\left\|\mathbf{y}\right\|_{2}}\zeta$, then $\mathbf{x}\in\mathcal{B}(R)$ and $\left\|\mathbf{x}-\mathbf{x}^{\ast}\right\|_{2}\leq\zeta$, where $\mathbf{x}$ is constructed from $\widetilde{\mu}$ as
Proof of Lemma˜11.
Eq.˜43 immediately implies $\mathbf{x}\in\mathcal{B}(R)$.
By Lemma˜7, we have $\left\|\mathbf{x}-\mathbf{x}^{\ast}\right\|_{2}\leq\left\|\widetilde{\mathbf{x}}-\mathbf{x}^{\ast}\right\|_{2}$. Then it suffices to prove $\left\|\widetilde{\mathbf{x}}-\mathbf{x}^{\ast}\right\|_{2}\leq\zeta$:
$\leq\sum_{i=1}^{d}v_{i}^{2}\left(1-\frac{1}{1+\frac{\left|\widetilde{\mu}-\mu^{\ast}\right|}{\lambda_{i}}}\right)^{2}\leq\sum_{i=1}^{d}v_{i}^{2}\left(\frac{\widetilde{\mu}-\mu^{\ast}}{\lambda_{i}}\right)^{2}\leq\left\|\mathbf{y}\right\|_{2}^{2}\left(\frac{\widetilde{\mu}-\mu^{\ast}}{\lambda_{d}}\right)^{2}<\zeta^{2}$
The first inequality is because $\lambda_{i}>0$, $\widetilde{\mu}\geq 0$ and $\mu^{\ast}\geq 0$; The second inequality employs the following inequality, $(1-\frac{1}{1+a})^{2}=\frac{a^{2}}{(a+1)^{2}}\leq a^{2}$ for any $a\geq 0$; The third inequality is uses $\sum_{i=1}^{d}v_{i}^{2}=\left\|\mathbf{y}\right\|_{2}^{2}$; The last inequality substitutes $|\widetilde{\mu}-\mu^{\ast}|<\frac{\lambda_{d}}{\left\|\mathbf{y}\right\|_{2}}\zeta$. ∎
Proof of Theorem˜3.
First, we justify the bisection. By Lemma˜10, the wanted zero satisfies $\mu^{\ast}\in((\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{d},(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1)\lambda_{1})=(a_{1},b_{1})$, which implies the initial interval of bisection (Line 3 in Algorithm˜4). Furthermore, $\rho$ monotonically decreases on $0,\infty)$, which implies the selection of $a_{t+1}$ and $b_{t+1}$ (Lines 4–6 in Algorithm˜4).
Then by Lemma˜11, Algorithm˜4 achieves an error $\left\|\mathbf{x}-\mathbf{x}^{\ast}\right\|_{2}\leq\zeta$.
Finally, we bound the number of arithmetic operations. Note that choice 1 and choice 2 are equivalent as
With choice 1, each iteration requires $d^{\omega}$ arithmetic operations, where the computational bottleneck lies in the matrix inversion $(A+\mu I)^{-1}$, resulting in a runtime of $O(d^{\omega}n)$. With choice 2, the tridiagonalization of $A$ requires $O(d^{3})$ arithmetic operations [Golub and Van Loan, 2013], and each iteration can solve $(C+\mu I)^{-1}\mathbf{w}$ using the Thomas algorithm with only $O(d)$ arithmetic operations [Golub and Van Loan, 2013], resulting in a runtime of $O(d^{3}+dn)$. ∎
Appendix C Proofs for Section˜4
In Section˜C.1, we prove Theorem˜4 based on Corollary 2 of Mehta [2017]. In Section˜C.2, we provide proofs for the ERM-based SXO methods discussed in Section˜4.3.
C.1 Proof of Theorem˜4

### v2

@@ C.1 Proof of Lemma˜2
Let $\Phi_{T}$ denote the sum of squared norms of updates, as in the left-hand side of Eq.˜14 in Lemma˜6. Note that $A_{t}=\epsilon I+\sum_{i=1}^{t}\nabla f_{t}(\mathbf{x}_{t})\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}$ in Algorithm˜1 and $\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}\leq G$ in Assumption˜2. Then by Lemma˜6 we have
@@
We prove Lemma˜2 by bounding $N$ with $\Phi_{T}$. Let $\{\tau_{i}\}_{i=1}^{N}\subseteq[T]$ denote all the rounds when LightONS.Core performs the Mahalanobis projection. With $\tau_{0}=1$, we have that for any $i\in[N]$,
@@
We remark that Lemma˜2 is tight in terms of $T$ up to logarithmic factors. Consider $d=1$ and $\nabla f_{t}(\mathbf{x}_{t})=1/\sqrt{T}$, then $A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})=\Omega(1/\sqrt{T})$ and $\|\sum_{t=1}^{T}A_{t}^{-1}\nabla f_{t}(\mathbf{x}_{t})\|_{2}=\Omega(\sqrt{T})$. We notice that Theorem 2.3 of Simchowit [2020] corroborates this point.
C.2 Proof of Theorem˜1
@@
The proof of Theorem˜1 follows as a specialization of the proof of Theorem˜2 given in Section˜C.4. The only substantive difference is that, in the present setting, the analysis is carried out directly on the original loss functions $f_{t}$, rather than on surrogate losses $g_{t}$.
@@
  • Choice of the expanded curvature parameter $\gamma$ in LightONS.Core. The selection of $\gamma=\frac{1}{2}\min\{\frac{2}{(k+1)DG},\alpha\}$ directly follows from Lemma˜1.
  • Regret analysis with selective projections. The regret analysis follows the same structure as Lemma˜11, which accounts for projections triggered only when the iterate exits the expanded domain $\widetilde{\mathcal{X}}_{k}$. Details are given in Section˜C.4.
  • Runtime analysis with selective projections. The runtime bound is derived in the same manner as in the proof of Theorem˜2 in Section˜C.4.
C.3 Proof of Lemma˜4
@@
By the selection of curvature parameters $\gamma_{0}$ (in Lemma˜1) and $\gamma^{\prime}$ (in Algorithm˜3), it holds that $\gamma^{\prime}\leq\gamma_{0}$. Thus by relaxing Eq.˜2 in Lemma˜1 we have
@@
$\leq\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})-\frac{\gamma^{\prime}}{2}\left(\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}(\mathbf{x}_{t}-\mathbf{u})\right)^{2}.$
To finish the proof, it suffices to show that, with $U(a)=a-\frac{\gamma^{\prime}}{2}a^{2}$,
@@
The preceding inequality holds because $U$ monotonically increases on $(-\infty,\frac{1}{\gamma^{\prime}}]$ and that
@@
The first inequality follows from Condition˜1; The second inequality from the Cauchy-Schwarz inequality $\mathbf{u}^{\mathsf{T}}\mathbf{v}\leq\left\|\mathbf{u}\right\|_{2}\left\|\mathbf{v}\right\|_{2}$; The last inequality from the selection of $\gamma^{\prime}$. ∎
C.4 Proof of Theorem˜2
@@
Based on Lemma˜10, we can decompose the regret of LightONS as the following lemma shows.
@@
Proof of Lemma˜11.
When $\left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}\leq kD/2$, no Mahalanobis projection is triggered, i.e., $\mathbf{y}_{t+1}=\widehat{\mathbf{y}}_{t+1}=\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})$, then for any $\mathbf{u}\in\mathbb{R}^{d}$, we have
@@
Otherwise, when $\left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}>kD/2$ and the Mahalanobis projection is triggered, by Lemma˜10, for any $\mathbf{u}\in\mathcal{X}$, we have
@@
$\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{A_{t}}^{2}\leq\left\|\left(\mathbf{y}_{t}-\mathbf{u}\right)-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})\right\|_{A_{t}}^{2}.$
Rearranging the preceding inequality yields the desired result of Eq.˜20. ∎
We note that Lemma˜11 ignores the truncation error of the Mahalanobis projections. The matrix factorization underlying Mahalanobis projections is related to eigendecomposition, equivalent to finding roots of a $d$-degree polynomial, which is not exactly solvable by finitely many arithmetic operations when $d\geq 5$. The next lemma complements the analysis. Thanks to FastProj (Algorithm˜4 in Section˜B.3), the truncation error only incurs a negligible additive $O(1/t^{2})$ term in the regret decomposition of Lemma˜11, which can be safely ignored in the final regret bound of LightONS.
@@
Let the Mahalanobis projection of Algorithm˜3 be implemented with Algorithm˜4 with $A=A_{t}$, $R=D/2$, $\mathbf{u}=\widehat{\mathbf{y}}_{t+1}$, $\zeta=\zeta_{t}$, $\underline{\lambda}=\epsilon$, $\overline{\lambda}=\epsilon+G^{2}t$, where $\zeta_{t}$ is defined as
@@
Proof of Lemma˜12.
It suffices to prove the case when $\left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}>kD/2$. Let $\mathbf{y}^{\ast}_{t+1}=\Pi_{\mathcal{B}(D/2)}^{A_{t}}\left[\widehat{\mathbf{y}}_{t+1}\right]$ be the exact Mahalanobis projection without truncation error, and $\bm{\delta}_{t}=\mathbf{y}^{\ast}_{t+1}-\mathbf{y}_{t+1}$ be the truncation error. By Lemma˜9, $\left\|\bm{\delta}_{t}\right\|_{2}=\left\|\mathbf{y}_{t+1}-\mathbf{y}^{\ast}_{t+1}\right\|_{2}\leq\zeta_{t}$; By Lemma˜10, for any $\mathbf{u}\in\mathcal{X}$,
@@
$\leq\left\|\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}A_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})-\mathbf{u}\right\|_{A_{t}}^{2}+\frac{2\gamma^{\prime}}{t^{2}}.$
The first inequality considers the operator norm of $A_{t}$ and uses that $\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{2}\leq c_{g}G$ (Condition˜1); The second inequality comes from the selection of $\zeta_{t}$ in Eq.˜21. ∎
@@
Plugging Lemma˜11 into Lemma˜4 yields the following inequalities:
@@
[…]
Telescoping the right-hand side establishes the desired regret of Eq.˜10a:
@@
Finally, the desired runtime of Eq.˜10b follows from the following two parts:
@@
  • Runtime aside from FastProj. In each round, the domain conversion of Lemma˜3 takes $O(\texttt{EP}_{\mathcal{X}}+d)$ time, updating and inverting $A_{t}$ as Eq.˜16 takes $O(d^{2})$ time, and other operations, such as computing the surrogate gradient as Eq.˜8b, take only $O(d)$ time. The overall runtime aside from FastProj is $O((\texttt{EP}_{\mathcal{X}}+d^{2})T)$.
@@
  • Runtime of FastProj. By Lemma˜7, the number of bisections $n_{t}=O(\log t)=O(\log T)$ and each bisection takes $O(d^{\omega})$ time with choice 1. Since LightONS.Core is a subroutine in LightONS, the number of calls to FastProj is at most $O((k-1)^{-1}d^{0.5}\sqrt{T/\epsilon})$ by Lemma˜2. The overall runtime of FastProj is $O((k-1)^{-1}d^{0.5}\sqrt{T/\epsilon}\cdot d^{\omega}\log T)$.
Appendix D Proofs for Section˜4
In this section, we provide the proofs and details for Section˜4.
D.1 Proof of Theorem˜3
Before proving Theorem˜3, we first give a more detailed presentation of the convergence rates induced by LightONS. With $T=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}\right)$, we have
$F(\bar{\mathbf{x}}_{T})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\leq\frac{1}{T}\left(\textsc{Reg}_{T}+4\sqrt{\frac{\textsc{Reg}_{T}}{2\gamma_{0}}\log\frac{4\log T}{\delta}}+\frac{8}{\gamma_{0}}\log\frac{4\log T}{\delta}\right)=O(\epsilon),$ (23a)
$\textsc{Runtime}\leq O\left(\frac{d^{3}}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}+\frac{d^{3.5}}{\sqrt{\epsilon}}\log\frac{d}{\epsilon}\log\frac{1}{\delta}\right)=\widetilde{O}\left(\frac{d^{3}}{\epsilon}\right),$ (23b)
where $\textsc{Reg}_{T}$ follows Eq.˜11a in Corollary˜1. With $\frac{1}{\delta}=T^{\prime}=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\right)$, we have
$\mathop{\mathbb{E}}\left[F(\bar{\mathbf{x}}_{T^{\prime}})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\right]\leq O(\epsilon),$ (24a)
$\textsc{Runtime}\leq O\left(\frac{d^{3}}{\epsilon}\log\frac{d}{\epsilon}+\frac{d^{3.5}}{\sqrt{\epsilon}}\log\frac{d}{\epsilon}\right)=\widetilde{O}\left(\frac{d^{3}}{\epsilon}\right).$ (24b)